A cylindrical container has one end opened and the other end closed. It has a circular base of radius r cm. Given that the total surface area of the container is 200π cm2.


(a) Show that the volume of the container is V = 100πr - πr^3/2 cm3

(b) Find the maximum value of V.


1
Expert's answer
2022-03-16T02:22:35-0400

Let height of the cylindrical container be h cm

So πr² + 2πrh = 200π

=> r² + 2rh = 200

=> h = "\frac{200-r²}{2r}"

So volume of the cylindrical container is , V =

"πr²h=(\frac{200-r²}{2r}) = πr(100-\frac{r²}{2})"

= "(100πr-\frac{πr³}{2})" cm³

Now we find maximum volume....

"\frac{dV}{dr} = 100π - \frac{3πr²}{2}"

"\frac{dV}{dr} = 0 =>" "100π - \frac{3πr²}{2}=0"

=> r² = "\frac{200}{3}"

=> r = "\frac{10\sqrt2}{\sqrt3}"

Now "\frac{d²V}{dr²} = - 3πr"

"[\frac{d²V}{dr²} ]_{r=\frac{10\sqrt2}{\sqrt3}}= - 3π*\frac{10\sqrt2}{\sqrt3} <0"


So volume is maximum when r= "\frac{10\sqrt2}{\sqrt3}"

So maximum volume is "(100πr-\frac{πr³}{2})"

= "\frac{π}{2}[200-r²]r"

= "\frac{π}{2}[200-\frac{200}{3}]\frac{10\sqrt2}{\sqrt3}"

= "\frac{π}{2}[\frac{400}{3}]\frac{10\sqrt2}{\sqrt3}"

= 1710.06644 cm³

= 1710.07 cm³ ( Correct up to two decimal places)